# Yongyang Cai  Khyati Malik   Hyeseon Shin

### v2

@@ 1 Introduction
While there is no general agreement among economists on whether carbon tax or ETS is better (Stavins, 2022), the literature have often considered two pricing instruments (trading versus taxes) as policy substitutes. For instance, some studies support carbon tax over ETS, as ETS is subject to price volatility (Nordhaus, 2007). Conversely, other studies advocate ETS over tax because allocation of permits in ETS allows more flexibility, and ETS faces less uncertainty in controlling the cumulative amount of carbon emissions compared to taxes (Keohane, 2009). ETS could be a preferred instrument over tax in practice, as free allowances could be negotiated among participating agents to redistribute burden and, thus, be used to gain political support (Harstad and Eskeland, 2010; Hahn and Stavins, 2011; Stavins, 2022). Newell and Pizer (2003) show that ETS is preferable over carbon tax in the long run, and Knopf et al. (2012) favor a global ETS because carbon taxes that increase over time may accelerate extraction of fossil fuels due to profit maximization of fossil fuel firms, an argument that has been made by other researchers as well (Sinn, 2008). Harstad and Eskeland (2010) argue that an ETS in a perfect market is the first-best system, but frequent government interventions to redistribute allocations among firms results in distortions in the market allocation. In a global ETS, emission permits will be traded globally with one price for all nations based on market forces, which may alleviate the problem of carbon leakage (Fowlie et al., 2016). For a comprehensive comparison on carbon tax and trading regimes, see e.g., Strand (2013); Schmalensee and Stavins (2017); Cai (2021); Stavins (2022). This study contributes to the literature by demonstrating that the two pricing instruments are not competing policies but rather complementary. A global ETS necessitates regional carbon taxes.
@@ 2 A Dynamic Regional Model of Climate and the Economy with an ETS
This section introduces a dynamic regional model of climate and the economy that integrates a global ETS between 12 aggregated regions. Our model framework extends the RICE model (Nordhaus and Yang, 1996; Nordhaus, 2010) by incorporating the global ETS in a dynamic setting and the TCRE climate system with annual time steps. The main objective of this study is to present a potential dynamic path of carbon emission prices in a noncooperative environment where each region aims to maximize its own social welfare by choosing the optimal amount of emission abatement, emission permits purchased from the global permits market, and the amount of its consumption. The model is based on the assumption that future regional emissions are strictly constrained by the emission caps set by the commitments under the Paris Agreement (and revised under the Glasgow Pact) and net zero emissions targets. In this study, we consider the case without uncertainty and without transaction cost in the global ETS.
The macroeconomic framework of our model employs a multi-regional representation of the Ramsey growth model. We follow the RICE model and assume that the world consists of 12 aggregated regions: the United States (US), the EU, Japan, Russia, Eurasia, China, India, Middle East (MidEast), Africa, Latin America (LatAm), Other High-Income countries (OHI) and other non-OECD Asia (OthAs). [^4] See Appendix 1 A.2 for the full list of countries and regional aggregation. Each region is indexed by $i\in\mathcal{I}$, where $\mathcal{I}$ is the set of regions. Time is discrete, with annual time steps indexed by $t=0,1,2,\cdots$, where the initial time period is the year 2020. All regions are forward-looking with complete information.
@@ 2.1 Emissions
The total industrial carbon emissions before abatement, in gigatonnes of carbon (GtC), is assumed to be proportional to the gross output for region $i$ at time $t$:
$E_{i,t}^{\mathrm{Total}}=\sigma_{i,t}Q_{i,t},$ (1)
where $Q_{i,t}$ is the gross output, and $\sigma_{i,t}$ is the exogenous carbon intensity, calibrated from the projections of GDP and emissions in Ueckerdt et al. (2019). We follow DICE and assume that the amount of emissions abated by each region, $E_{i,t}^{A}$, is expressed as a fraction of the total industrial emissions, with an emission control rate $\mu_{i,t}\in[0,1]$. [^5] Instead of an emission control rate, researchers also model abatement through reduced use of fossil fuel energy inputs in a production function (see, e.g., Bosetti et al., 2006; Bauer et al., 2012; Golosov et al., 2014; Baldwin et al., 2020). However, these models require disaggregation of fossil fuel energy firms, renewable energy firms, final-goods producing firms, and other sectors, so they are often complicated. The emission control rate concept simplifies this. That is,
@@
Thus the net emissions, $E_{i,t}$, becomes
$E_{i,t}=E_{i,t}^{\mathrm{Total}}-E_{i,t}^{A}.$ (3)
We consider a global ETS, or a cap-and-trade system, in which each region is provided with an emission allowance and can trade emission permits with other regions. Regions with net emissions exceeding the emission caps can purchase emission permits from other regions. Regions with emissions below their allowance can sell their emission permits to others. The emission cap constraint is represented by
@@
where $\overline{E}_{i,t}$ is the emission cap assigned to region $i$ at time $t$, and $E_{i,t}^{P}$ denotes the amount of emissions traded with other regions. Note that $E_{i,t}^{P}>0$ indicates that region $i$ is a net buyer of emission permits at time $t$, while $E_{i,t}^{P}<0$ implies region $i$ is a net seller of emission permits at time $t$. In a model that does not consider emission permit trade, we simply apply $E_{i,t}^{P}=0$ for all $i$ and $t$. The Paris Agreement establishes upper limits for net emissions to ensure that the increase in global temperature remains well below 2 degrees Celsius above preindustrial levels by the end of this century.
@@ 2.2 Climate System
where the cumulative emissions at 2020, $\mathcal{E}_{0}$, is chosen such that the initial global mean temperature $T_{0}$ is 1.2 degrees Celsius above the pre-industrial level. Here $\zeta=0.0021$ is the contribution rate of CO2 equivalent to temperature, calibrated with the projections of emissions and temperatures in the four Representative Carbon Concentration Pathways (RCPs): RCP 2.6, RCP 4.5, RCP 6.0, and RCP 8.5 (Meinshausen et al., 2011). Figure A.1 in Appendix A.3 shows that the calibrated TCRE climate system can match all four RCP temperature pathways using their associated RCP emission pathways. Cumulative global emissions evolve according to

### v3

@@ 1 Introduction
While there is no general agreement among economists on whether a carbon tax or an ETS is better (Stavins,, 2022), the literature have often considered these two pricing instruments (trading versus taxes) as policy substitutes. For instance, some studies support carbon tax over ETS, highlighting concerns about price volatility in ETS (Nordhaus,, 2007), or the presence of uncertainty regarding emission abatement costs (Newell and Pizer,, 2003). For instance, Newell and Pizer, (2003) find that carbon taxes can yield higher welfare benefits than an ETS under such uncertainties. Conversely, other studies advocate ETS over carbon tax because allocation of permits in ETS allows for more flexibility, and ETS faces less uncertainty in controlling the cumulative amount of carbon emissions compared to taxes (Keohane,, 2009). Harstad and Eskeland, (2010), Hahn and Stavins, (2011), and Stavins, (2022) argue that from a practical perspective, ETS may be a preferred instrument over tax, as free allowances could be negotiated among participating agents to redistribute burden and, thus be used to gain political support. Knopf et al., (2012) favor a global ETS because carbon taxes that increase over time may accelerate extraction of fossil fuels due to profit maximization of fossil fuel firms, an argument that has been made by other researchers as well (Sinn,, 2008). In a global ETS, emission permits will be traded globally with one price for all nations based on market forces, which may alleviate the problem of carbon leakage (Fowlie et al.,, 2016). However, Harstad and Eskeland, (2010) argue that although an ETS in a perfect market is the first-best system, frequent government interventions to redistribute allocations among firms results in distortions in the market allocation. For a comprehensive comparison on carbon tax and trading regimes, see e.g., Strand, (2013); Schmalensee and Stavins, (2017); Cai, (2021); Stavins, (2022). This study contributes to the literature by demonstrating that the two pricing instruments are not competing policies but rather complementary.
2 Optimal Carbon Tax under the ETS
While carbon taxes and ETS are often considered as substitutable policies, we demonstrate that each country needs to implement an optimal carbon tax as a complementary policy within a global ETS regime to achieve optimal level of emissions. Before introducing the full multi-region dynamic general equilibrium model, it is helpful to first explore the optimal carbon tax within a static framework. Under the global ETS framework, we assume that each region $i$ is allocated with an exogenous level of emission cap ($\bar{E}_{i}$), which is the upper limit of emissions allowed for that region and is strictly enforced. The region’s gross emissions before abatement is $E_{i}^{\mathrm{Gross}}$, its emissions net of abatement is $E_{i}$, the amount of emission permits purchased (or sold) in the market is $E_{i}^{P}$ (if $E_{i}^{P}<0$, it implies that the region sells emission permits) so we have the emission cap constraint $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$, where the Nash equilibrium permit price is $m$ under the market clearing condition $\sum_{j\in\mathcal{I}}E_{j}^{P}=0$. Here we assume that each region is a price taker. The region’s marginal abatement cost (MAC) function, denoted $\mathrm{MAC}_{i}(E_{i})$, is assumed to be monotonically decreasing over emissions $E_{i}$ with $\mathrm{MAC}_{i}(E_{i}^{\mathrm{Gross}})=0$. A social planner of the region solves the following minimization problem
[…]
where the first term of the objective function is the regional damage from the global emissions $\sum_{j\in\mathcal{I}}E_{j}$ (which is the present value of future damage from the global emissions), the second term is the cost from purchasing permits at the price $m$ (or the profit from selling permits), and the third term is the total abatement cost for reducing emissions by the amount $(E_{i}^{\mathrm{Gross}}-E_{i})$. For simplicity, we assume that emissions cannot be negative in our model.
In the static Nash equilibrium of all regional social planners, every region solves equation (1) simultaneously with the market clearing condition $\sum_{j\in\mathcal{I}}E_{j}^{P}=0$, which implies $\sum_{j\in\mathcal{I}}E_{j}\leq\sum_{j\in\mathcal{I}}\overline{E}_{j}$. Let $E_{j}^{*}$ be the Nash equilibrium emissions for each region $j$ and let $m^{*}$ be the Nash equilibrium price. Thus, from the Karush-Kuhn-Tucker conditions of (1), the solution to the region’s optimal emissions leads to the following relationship between the marginal regional damage, the MAC, and the permit price:
$D_{i}^{\prime}\left(\sum_{j\in\mathcal{I}}E_{j}^{*}\right)=\mathrm{MAC}_{i}\left(E_{i}^{*}\right)-m^{*},$ (2)
when $E_{i}^{*}>0$. The marginal regional damage can also be regarded as the regional social cost of carbon (SCC). Note that in the static Nash equilibrium with $E_{i}^{*}>0$, the permit price $m^{*}$ should be equal to the shadow price of the inequality $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$. That is, if the inequality $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$ is not binding, or $\sum_{j\in\mathcal{I}}E_{j}^{*}<\sum_{j\in\mathcal{I}}\overline{E}_{j}$ (i.e., there is an over-supply of emission permits), then the permit price is zero. Alternatively, if the permit price is nonzero, then the inequality $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$ must be binding, so the supply of emission permits is equal to the demand (i.e., $\sum_{j\in\mathcal{I}}E_{j}^{*}=\sum_{j\in\mathcal{I}}\overline{E}_{j}$). We emphasize that the assumption commonly used in the literature—that the MAC equals the permit price (e.g., Keohane,, 2009; Nordhaus, 2010b, )—however, does not hold in this case, unless the regional marginal damage or SCC is zero, according to equation (2). Moreover, we also numerically verify in our dynamic Nash equilibrium that, when $E_{i}^{*}>0$, the regional MAC is also strictly larger than the permit price, and the difference is the regional SCC.
An important issue arises when a regional social planner cannot directly enforce its optimal emissions level on firms, the actual emission producers in the economy. To address this, let us assume that the social planner can employ carbon taxation as a policy instrument to resolve the discrepancy in optimal emissions between the firms and the regional social planner. Consider a representative firm in region $i$ that produces emissions and operates under the region’s MAC function and emissions cap constraint in a static framework. Since the firm does not internalize the damage from emissions, we assume that the social planner of region $i$ levies a tax on the firm of $\tau_{i}$ per unit of emissions. The firm solves the following minimization problem:
[…]
where the first term of the objective function is the total tax on emissions. In the static Nash equilibrium of all representative firms, every firm solves equation (3) simultaneously with the market clearing condition $\sum_{j\in\mathcal{I}}E_{j}^{P}=0$. Let $E_{j}^{*,\mathrm{Firm}}$ be the Nash equilibrium emissions for each firm $j$ and let $m^{*,\mathrm{Firm}}$ be the Nash equilibrium price. Thus, from the Karush-Kuhn-Tucker conditions of (3), the solution to the optimal emissions of the firm leads to the following relationship between the carbon tax, the MAC, and the permit price:
$\tau_{i}=\mathrm{MAC}_{i}(E_{i}^{*,\mathrm{Firm}})-m^{*,\mathrm{Firm}},$ (4)
when $E_{i}^{*,\mathrm{Firm}}>0$. If the regional social planner imposes the tax to be equal to the regional SCC, i.e., the regional marginal damage, then from equations (2) and (4), we have $E_{i}^{*,\mathrm{Firm}}=E_{i}^{*}$ and $m^{*,\mathrm{Firm}}=m$. In other words, the optimal regional carbon tax is endogenously determined by the gap between the MAC at the regional social planner’s optimal emissions and the market price of emission permits, capturing the idea that the regional social planner can choose the level of carbon tax such that the firm can internalize the regional SCC and achieve its emissions level of $E_{i}^{*,\mathrm{Firm}}=E_{i}^{*}$. When $E_{i}^{*,\mathrm{Firm}}=0$, the Karush-Kuhn-Tucker conditions of (3) shows that
$\tau_{i}\geq\mathrm{MAC}_{i}(E_{i}^{*,\mathrm{Firm}})-m^{*,\mathrm{Firm}}.$ (5)
That is, the carbon tax could be larger than the gap between the MAC of a region and the market price of emission permits. In particular, the regional $\mathrm{MAC}_{i}(E_{i}^{*,\mathrm{Firm}})$ at $E_{i}^{*,\mathrm{Firm}}=0$ could be smaller than the global permit price $m^{*,\mathrm{Firm}}$, making the right hand side of the inequality (5) to be negative, so the optimal carbon tax should be zero or the gap $\mathrm{MAC}_{i}(E_{i}^{*})-m^{*}$ if it is positive.
Therefore, for any regional social planner’s optimal emissions $E_{i}^{*}\geq 0$, we have the following general formula to choose the optimal carbon tax:
$\tau_{i}=\max\{0,\ \mathrm{MAC}_{i}(E_{i}^{*})-m^{*}\},$ (6)
such that the firm’s optimal emissions are consistent with the social planner’s (i.e., $E_{i}^{*,\mathrm{Firm}}=E_{i}^{*}$). The simple relationship between the optimal carbon tax, the MAC, and the permit price is also verified in our numerical solutions of the dynamic regional model. Note that this tax is contingent on the imposition of the emission cap on the firm, and the tax level itself cannot make the firm’s emissions to be $E_{i}^{*}$ unless $m^{*}=0$.
It can be easily shown that our argument is consistent with the existing literature when there is no ETS. The cost minimization problem of the representative firm facing the emission cap regulation without the ETS becomes:
[…]
From the Karush-Kuhn-Tucker conditions of (7), when the firm’s optimal solution $E_{i}^{*,\mathrm{Firm}}=E_{i}^{*}$ is not binding at its lower or upper bound, we have the well known equality:
$\tau_{i}=\mathrm{MAC}_{i}(E_{i}^{*}).$ (8)
This implies that the optimal carbon tax should be set equal to the MAC at the regional social planner’s optimal emissions $E_{i}^{*}$, provided $E_{i}^{*}$ is not binding. This result has been widely used in the literature (e.g., Nordhaus,, 2014; Cai et al.,, 2017; Cai and Lontzek,, 2019). Note that our formula (6) under an ETS with zero permit price is consistent with (8) under no ETS.
A regional social planner in each region $i$ has two targets: one is its exogenous NDC commitment $\overline{E}_{i}$, and the other is the regional social planner’s endogenous optimal emissions $E_{i}^{*}$ after permit trading (note that $\overline{E}_{i}$ and $E_{i}^{*}$ are unequal in general [^5] We believe this is a realistic consideration, as emission cap pathways may not always align precisely with regional optimal emissions over time. This discrepancy arises because the negotiation of emission permit allocations is often time-consuming and influenced by political considerations, which may not lead to economically optimal outcomes. Our model simulation shows that the emission cap pathways constructed from the Paris Agreement and the Glasgow Climate Pact indeed deviate from the optimal regional emissions.) in the Nash equilibrium of all 12 regional social planners. That is, the region $i$’s emissions should be $E_{i}^{*}$, which should also satisfy $E_{i}^{*}\leq\overline{E}_{i}+E_{i}^{P}$ after permit trading in the Nash equilibrium. Several policy instruments can be applied to enforce the firms to meet the targets, assuming that the firms can buy or sell emision permits. One such instrument is a quantity policy, in which the regional social planner distributes $\overline{E}_{i}$ (e.g., by means of an auction or grandfathering) to its firms, ensuring the firms’ aggregate emissions, $E_{i}^{*,\mathrm{Firm}}$, do not exceed the exogenous allocated cap after permit trading (i.e., $E_{i}^{*,\mathrm{Firm}}\leq\overline{E}_{i}+E_{i}^{P}$) in the Nash equilibrium of firms of all regions. [^6] The representative firm framework can be generalized to an environment with multiple firms. Assume that there are $n$ homogenous firms in the region $i$, each having $\overline{E}_{i}/n$ as their caps, with a firm’s MAC function at its emissions $e$ is defined as $\mathrm{MAC}_{i}(ne)$, where $\mathrm{MAC}_{i}(\cdot)$ is the aggregate regional MAC function used by the regional social planner and $E=ne$ is the aggregate regional emissions. Under this definition, the aggregate total abatement cost of all firms in region $i$ is equal to the integral of $\mathrm{MAC}_{i}(E)$ on $[E_{i},E_{i}^{\mathrm{Gross}}]$, when each firm’s emission is $E_{i}/n$ while $E_{i}^{\mathrm{Gross}}/n$ is each firm’s upper bound under no abatement. In this case, the Nash equilibrium of the economy with multiple firms is identical to the equilibrium with a representative firm. However, under a quantity-only control policy, $E_{i}^{*,\mathrm{Firm}}$ is not equal to $E_{i}^{*}$, as the zero-tax Nash equilibrium of the firms, $\left\{m^{*,\mathrm{Firm}},E_{i}^{*,\mathrm{Firm}}:j\in\mathcal{I}\right\}$, is not the Nash equilibirum of the regional social planners, $\left\{m^{*},E_{j}^{*}:j\in\mathcal{I}\right\}$. This becomes particularly evident in the following extreme case: assume that $\overline{E}_{i}$ is larger than $E_{i}^{\mathrm{Gross}}$ for every region, which leads to $m^{*}=0$ under the ETS. The firms’ Nash equilibrium without a tax yields $E_{i}^{*,\mathrm{Firm}}=E_{i}^{\mathrm{Gross}}$, with $\mathrm{MAC}_{i}(E_{i}^{*,\mathrm{Firm}})=m^{*,\mathrm{Firm}}=0$. In contrast, the regional social planners’ Nash equilibrium leads to $E_{i}^{*}<E_{i}^{\mathrm{Gross}}$, with $\mathrm{MAC}_{i}(E_{i}^{*})>m^{*}=0$ in general, similar to a model without emission caps. Note that this quantity policy cannot directly distribute $E_{i}^{*}$, as doing so would either result in the region violating its commitment if $E_{i}^{*}>\overline{E}_{i}$ or losing the benefit in selling permits if $E_{i}^{*}<\overline{E}_{i}$. Another instrument is a carbon tax policy, in which the region $i$ imposes a carbon tax $\tau_{i}$ on the firms, e.g., $\tau_{i}=\mathrm{MAC}_{i}(\overline{E}_{i})$ or $\tau_{i}=\mathrm{MAC}_{i}(E_{i}^{*})$. However, a carbon tax policy alone can target only one emission level, either $\overline{E}_{i}$ or $E_{i}^{*}$, in each region, while $\overline{E}_{i}\neq E_{i}^{*}$ in general. Therefore, neither instrument can achieve both targets simultaneously.
We argue that in the presence of a global ETS, it is optimal for each region $i$ to impose both the quantity policy and an additional carbon tax policy simultaneously on firms to enforce both targets. Here the additional carbon tax for region $i$ is chosen to be $\tau_{i}=\mathrm{MAC}_{i}(E_{i}^{*})-m^{*}$ when $E_{i}^{*}>0$, under the Nash equilibirum of the planners, $\left\{m^{*},E_{j}^{*}:j\in\mathcal{I}\right\}$. If some regions impose only the quantity policy while the others impose both policies, this leads to a Nash equilibrium of firms, which is not the Nash equilibirum of the planners. Note that a tax has no impact on the regional social planners’ decisions; $E_{i}^{*}$ remains the optimal decisions of the regional social planners regardless of the tax levels. However, the tax can influence the firms’ decisions. Therefore, from the regional social planners’ perspective, any Nash equilibrium of firms deviating from $\left\{m^{*},E_{j}^{*}:j\in\mathcal{I}\right\}$ is worse, and the regions should always impose both the quantity policy and the optimal additional carbon taxes simultaneously on the firms to meet both targets. [^7] Our argument is based on the assumption that the emission caps are exogenous in the regional level.
@@ 3 A Dynamic Regional Model of Climate and the Economy with an ETS
We now introduce a dynamic regional model of climate and the economy that integrates a global ETS between 12 aggregated regions. Our model framework extends the RICE model (Nordhaus and Yang,, 1996; Nordhaus, 2010b, ) by incorporating the global ETS in a dynamic setting and the TCRE climate system with annual time steps. The main objective of this study is to present a potential dynamic path of carbon emission prices in a noncooperative environment where each region aims to maximize its own social welfare by choosing the optimal amount of emission abatement, emission permits purchased from the global permits market, and the amount of its consumption. The model is based on the assumption that future regional emissions are strictly constrained by the emission caps set by the commitments under the Paris Agreement (and revised under the Glasgow Pact) and net zero emissions targets. In this study, we consider the case without uncertainty and without transaction cost in the global ETS.
The macroeconomic framework of our model employs a multi-regional representation of the Ramsey growth model. We follow the RICE model and assume that the world consists of 12 aggregated regions: the United States (US), the EU, Japan, Russia, Eurasia, China, India, Middle East (MidEast), Africa, Latin America (LatAm), Other High-Income countries (OHI) and other non-OECD Asia (OthAs). [^8] See Appendix 1 A.2 for the full list of countries and regional aggregation. Each region is indexed by $i\in\mathcal{I}$, where $\mathcal{I}$ is the set of regions. Time is discrete, with annual time steps indexed by $t=0,1,2,\cdots$, where the initial year is the year 2020. All regions are forward-looking with complete information.
@@ 3.1 Emissions
The gross carbon emissions before abatement, in gigatonnes of carbon (GtC), is assumed to be proportional to the gross output for region $i$ at time $t$:
$E_{i}^{\mathrm{Gross}}=\sigma_{i,t}Q_{i,t},$ (9)
where $Q_{i,t}$ is the gross output, and $\sigma_{i,t}>0$ is the exogenous carbon intensity, calibrated from the projections of GDP and emissions in Ueckerdt et al., (2019). We follow DICE and assume that the amount of emissions abated by each region, $E_{i,t}^{A}$, is expressed as a fraction of the gross emissions, with an emission control rate $\mu_{i,t}\in[0,1]$. [^9] Instead of an emission control rate, researchers also model abatement through reduced use of fossil fuel energy inputs in a production function (see, e.g., Bosetti et al.,, 2006; Bauer et al.,, 2012; Golosov et al.,, 2014; Baldwin et al.,, 2020). However, these models require disaggregation of fossil fuel energy firms, renewable energy firms, final-goods producing firms, and other sectors, so they are often complicated. The emission control rate concept simplifies this. That is,
@@
Thus the emissions net of abatement, $E_{i,t}$, becomes
$E_{i,t}=E_{i}^{\mathrm{Gross}}-E_{i,t}^{A}.$ (11)
We consider a global ETS, or a cap-and-trade system, in which each region is provided with an emission allowance and can trade emission permits with other regions. Regions with emissions exceeding the emission caps can purchase emission permits from other regions. Regions with emissions below their allowance can sell their emission permits to others. The emission cap constraint is represented by
@@
where $\overline{E}_{i,t}$ is the emission cap assigned to region $i$ at time $t$, and $E_{i,t}^{P}$ denotes the amount of emissions traded with other regions. Note that $E_{i,t}^{P}>0$ indicates that region $i$ is a net buyer of emission permits at time $t$, while $E_{i,t}^{P}<0$ implies region $i$ is a net seller of emission permits at time $t$. In a model that does not consider emission permit trade, we simply apply $E_{i,t}^{P}=0$ for all $i$ and $t$. The Paris Agreement establishes upper limits for emissions to ensure that the increase in global temperature remains well below 2°C above pre-industrial levels by the end of this century.
@@ 3.2 Climate System
where the cumulative emissions at 2020, $\mathcal{E}_{0}$, is chosen such that the initial global mean temperature $T_{0}$ is 1.2°C above the pre-industrial level. Here $\zeta=0.0021$ is the contribution rate of CO2 equivalent to temperature, calibrated with the projections of emissions and temperatures in the four Representative Carbon Concentration Pathways (RCPs): RCP 2.6, RCP 4.5, RCP 6.0, and RCP 8.5 (Meinshausen et al.,, 2011). Figure A.1 in Appendix A.3.1 shows that the calibrated TCRE climate system can match all four RCP temperature pathways using their associated RCP emission pathways. Cumulative global emissions evolve according to